# Some geometric series for Euler’s constant

### 07 May 2026 v2

@@ 4. The asymptotic behavior of $(e_{m})$
The Bernoulli polynomials and numbers used in (13) are defined by $\sum_{n=0}^{\infty}B_{n}(x)\frac{t^{n}}{n!}=\frac{t\mathsf{e}^{xt}}{\mathsf{e}^{t}-1}$, $B_{n}(x)=\sum_{k=0}^{n}\binom{n}{k}B_{k}x^{n-k}$, $B_{0}=1$, $B_{1}=-\frac{1}{2}$, $B_{n}=B_{n}(0)$, $(m+1)\sum_{0\leq n<N}n^{m}=B_{m+1}(N)-B_{m+1}(0)$. Formula (13) has large contributions of both signs and does not look promising numerically. In the sequel we will only use (12).
@@
A variant of (13) is (see Equation (22) for the Euler polynomials):
$e_{m}$ $=\sum_{l=0}^{\infty}(l+1)\frac{B_{m+1}(2^{l+1})-2^{m+1}B_{m+1}(2^{l})+(2^{m+1}-1)B_{m+1}(0)}{2^{(l+1)(m+1)}}$
$=\frac{m+1}{2}\sum_{l=0}^{\infty}(l+1)\frac{E_{m}(2^{l+1})-E_{m}(0)}{2^{(l+1)(m+1)}}\,.$
One only needs to write the numerator as $(l+1)(v_{l+1}-2^{m+1}v_{l})$ with $v_{l}=B_{m+1}(2^{l})-B_{m+1}(0)-2^{(m+1)l}$ and rearrange the sum. Details are left to the reader.
We will derive in this section from Equation (12) some asymptotic for $e_{m}$ as $m\to\infty$. From (11), the $l$th term is non-negative and bounded above by $(m+1)2^{-l-1}$. Let $L_{m}=\lfloor 2\log_{2}(m+1)\rfloor$. The sum of the contributions for $l>L_{m}$ is bounded above by $(m+1)2^{-L_{m}-1}$, hence, by $(m+1)^{-1}$.
To approximate the ($l$ dependent) moments $p_{l}=2^{-l}\sum_{0\leq n<2^{l}}(2^{-l}n)^{m}$ for $l\leq L_{m}$, we shall apply the same technique as in [6, Prop. 1]. For the convenience of the reader, the presentation here is entirely self-contained. [^1] In v1 of [6], which is the current version at time of writing, the Proposition 1 is formulated in a way which does not make explicit the implied constants in the big-O, although this is quite easy to extract from the proof there. As here we need to apply this to mutliple probability measures, we had anyhow to reproduce parts of the argument from [6] in order to control the uniformity. Certainly, the published version of [6] will make more explicit the big-O bound. Let us define
$q_{l}=2^{-l}\sum_{0\leq n<2^{l}}\mathsf{e}^{-m(1-2^{-l}n)}=\frac{1-\mathsf{e}^{-m}}{2^{l}(\mathsf{e}^{m2^{-l}}-1)}\,.$ (14)
There holds $0\leq x\mathsf{e}^{-x}\leq\mathsf{e}^{-1}$ for $x\geq 0$, hence $x^{m}\leq\mathsf{e}^{-m(1-x)}$. In particular $(2^{-l}n)^{m}\leq\mathsf{e}^{-m(1-2^{-l}n)}$ and $0\leq p_{l}\leq q_{l}$. We split $q_{l}-p_{l}$ into two parts:
$q_{l}-p_{l}=2^{-l}(\sum_{0\leq n<2^{l-1}}+\sum_{2^{l-1}\leq n<2^{l}})\biggl(\mathsf{e}^{-m(1-2^{-l}n)}-(2^{-l}n)^{m}\biggr)$ (15)
We bound trivially the first part by
$2^{-l}\sum_{0\leq n<2^{l-1}}\mathsf{e}^{-m(1-2^{-l}n)}\leq\int_{0}^{\frac{1}{2}}\mathsf{e}^{m(x-1)}\mathrm{d}x=\frac{1}{m}\mathsf{e}^{-m/2}(1-\mathsf{e}^{-m/2}).$
This, and everything following applies also to non-integer positive $m$’s, the quantity $e_{m}$ can be defined for them using (12).
Regarding the second part in Equation (15), we use (as in the proof of [6, Prop. 1]) the bound
$m\geq 0,\;\frac{1}{2}\leq x\leq 1\implies\mathsf{e}^{-m(1-x)}-x^{m}\leq m(1-x)^{2}\mathsf{e}^{-m(1-x)}.$
So, for positive $m$:
$0\leq q_{l}-p_{l}\leq\frac{\mathsf{e}^{-m/2}}{m}+2^{-l}\sum_{2^{l-1}\leq n<2^{l}}m(1-2^{-l}n)^{2}\mathsf{e}^{-m(1-2^{-l}n)}$
(Note that if $l=0$, the sum above is empty and has value 0; so we now assume $l\geq 1$ and the bound we obtain also applies with $l=0$). Let $\psi_{m}(x)=\frac{\mathrm{d}}{\mathrm{d}x}\left((1-x)^{2}\mathsf{e}^{-m(1-x)}\right)=\bigl(m(1-x)-2\bigr)(1-x)\mathsf{e}^{-m(1-x)}$. We write:
$2^{-l}\sum_{2^{l-1}\leq n<2^{l}}m(1-2^{-l}n)^{2}\mathsf{e}^{-m(1-2^{-l}n)}=\frac{m}{8}\mathsf{e}^{-m/2}+m2^{-l}\sum_{2^{l-1}\leq n<2^{l}}\int_{\frac{1}{2}}^{2^{-l}n}\psi_{m}(y)\mathrm{d}y$
Using the probability measure $\mu_{l}=2^{-l}\sum_{0\leq n<2^{l}}\delta_{2^{-l}n}$, which has the property $\mu_{l}([y,1])\leq 1-y$ for every $y\in[0,1]$, we can rewrite the second summand on the right-hand side as:
$)\displaystyle m\int_{[\frac{1}{2},1\rparen}\int_{\frac{1}{2}}^{x}\psi_{m}(y)\mathrm{d}y\,\mu_{l}(\mathrm{d}x$) $\leq m\int_{\frac{1}{2}\leq y\leq x<1}\max(\psi_{m}(y),0)\mathrm{d}y\,\mu_{l}(\mathrm{d}x)$
$\leq m\int_{\frac{1}{2}}^{1}\max(\psi_{m}(y),0)(1-y)\mathrm{d}y$
$=m\int_{\frac{1}{2}}^{\max(1-\frac{2}{m},\frac{1}{2})}\bigl(m(1-y)-2\bigr)(1-y)^{2}\mathsf{e}^{-m(1-y)}\mathrm{d}y$
$=m^{-2}\int_{\min(2,m/2)}^{m/2}(t-2)t^{2}\mathsf{e}^{-t}\mathrm{d}t\leq 18\mathsf{e}^{-2}m^{-2}.$
We have obtained, for positive $m$ (with some loss of precision near zero, but we are only interested into $m\to\infty$):
$0\leq q_{l}-p_{l}\leq(m^{-1}+\frac{m}{8})\mathsf{e}^{-m/2}+\frac{18}{\mathsf{e}^{2}m^{2}}\leq C_{1}m^{-2},$ (16)
with $C_{1}=\max_{t\geq 0}(t+t^{3}/8)\mathsf{e}^{-t/2}+18\mathsf{e}^{-2}$. It turns out that the maximum is attained at $t\approx\numprint{5.04}$ and its value is $\approx\numprint{1.693}$ and $18\mathsf{e}^{-2}\approx\numprint{2.436}$ and $C_{1}<\numprint{4.13}$. Of course when $m$ is large $18\mathsf{e}^{-2}+\epsilon$ gives a tighter bound.
The proof applies to any positive $m$, and for $m$ a positive integer it becomes a statement involving the values of Bernoulli polynomials at $N=2^{l}$. That $N$ was a power of $2$ is not important. We formulate here a dedicated statement, as we could not find such property in the standard references.
@@
For any two positive integers $m$ and $n$, the following uniform estimate holds:
$0\leq\frac{1-\mathsf{e}^{-m}}{n\bigl(\mathsf{e}^{mn^{-1}}-1\bigr)}-\frac{B_{m+1}(n)-B_{m+1}(0)}{(m+1)n^{m+1}}\leq\frac{5}{m^{2}}.$ (17)

### 15 May 2026 v4

@@ 4. The coefficients $e_{m}$ via power sums
A variant of (16) is:
$e_{m}=\sum_{l=0}^{\infty}(l+1)\frac{B_{m+1}(2^{l+1})-2^{m+1}B_{m+1}(2^{l})+(2^{m+1}-1)B_{m+1}(0)}{2^{(l+1)(m+1)}}$ (17)
One only needs to write the numerator as $(l+1)(v_{l+1}-2^{m+1}v_{l})$ with $v_{l}=B_{m+1}(2^{l})-B_{m+1}(0)-2^{(m+1)l}$ and rearrange the sum. Details are left to the reader. In terms of the Euler polynomials (see Equation (32)), this is:
$\frac{m+1}{2}\sum_{l=0}^{\infty}(l+1)\frac{E_{m}(2^{l+1})-E_{m}(0)}{2^{(l+1)(m+1)}}\,.$
5. Approximations
Let us make the general definition, for $n$ a positive integer:
$p_{m}(n)=n^{-1}\sum_{0\leq k<n}(n^{-1}k)^{m}\,.$ (18)
They are the moments of the discrete probability measure $\sum_{0\leq k<n}n^{-1}\delta_{k/n}$. It is not needed here for $m$ to be an integer, and in what follows we let $e_{m}$ for $m$ real positive be given by Equation (15), i.e.:
$e_{m}=\sum_{l=0}^{\infty}\bigl(1-(m+1)p_{m}(2^{l})\bigr)\,$, (19)
and we study the asymptotic for real positive $m$ going to infinity.
There holds
$(m+1)p_{m}(n)=\frac{B_{m+1}(n)-B_{m+1}(0)}{n^{m+1}}\,.$ (20)
Let us start with the examination of (18) via “pointwise” approximations. We will use the same inequalities as in [6, Prop. 1]:
$0\leq x$ $\implies 0\leq\mathsf{e}^{-m(1-x)}-x^{m}\,$,
$\frac{1}{2}\leq x\leq 1$ $\implies 0\leq\mathsf{e}^{-m(1-x)}-x^{m}\leq m(1-x)^{2}\mathsf{e}^{-m(1-x)}\,.$ (21)
Inequality (21), with $x=1-n^{-1}k$, $1\leq k\leq\frac{n}{2}$, is found in a similar context also in [14, answer to 6.3-34]. It suggests approximating $p_{m}(n)$ by
$n^{-1}\sum_{0\leq k<n}\mathsf{e}^{-m(1-n^{-1}k)}=\frac{1-\mathsf{e}^{-m}}{n(\mathsf{e}^{n^{-1}m}-1)}\,$,
which we rewrite as $(1-\mathsf{e}^{-m})q_{m}(n)$, defining for this purpose:
$q_{m}(n)=\frac{1}{n(\mathsf{e}^{n^{-1}m}-1)}\,.$ (22)
But the approximation of $p_{m}(n)$ by $q_{m}(n)$ becomes poor for $n$ large, as:
$\lim_{n\to\infty}p_{m}(n)=\int_{0}^{1}x^{m}\mathrm{d}x=(m+1)^{-1},\text{ while }\lim_{n\to\infty}q_{m}(n)=m^{-1}.$
Thus, we should either switch to $q_{m+1}(n)$ (we shall see later that it arises very naturally from Equation (20)), or approximate $(m+1)p_{m}(n)$ via $mq_{m}(n)$.
We start with a Lemma which uses Equation (21) to compare $p_{m}(n)$ with $(1-\mathsf{e}^{-m})q_{m}(n)$ uniformly for the entire $n$ range.
Lemma 1.
There is a positive constant $K_{1}<3.1$ such that for every real positive $m$ and every positive integer $n$, there holds:
$0\leq(1-\mathsf{e}^{-m})q_{m}(n)-p_{m}(n)\leq\frac{K_{1}}{m^{2}}\,.$ (23)
Proof.
We write:
$(1-\mathsf{e}^{-m})q_{m}(n)-p_{m}(n)\begin{aligned} &=\Bigl(n^{-1}\sum_{0\leq k<n/2}+n^{-1}\sum_{n/2\leq k<n}\Bigr)\bigl(\mathsf{e}^{-m(1-n^{-1}k)}-(n^{-1}k)^{m}\bigr)\\ &=S_{1}+S_{2}\,.\end{aligned}$
We bound $S_{1}$ (trivially) for $n$ even by $\int_{0}^{\frac{1}{2}}\mathsf{e}^{-m(1-x)}\mathrm{d}x=m^{-1}(\mathsf{e}^{-m/2}-\mathsf{e}^{-m})$ and for $n$ odd by $\int_{0}^{\frac{1}{2}-(2n)^{-1}}\mathsf{e}^{-m(1-x)}\mathrm{d}x+2\int_{\frac{1}{2}-(2n)^{-1}}^{\frac{1}{2}}\mathsf{e}^{-m(1-x)}\mathrm{d}x=m^{-1}\bigl(\mathsf{e}^{-m/2}-\mathsf{e}^{-m}+\mathsf{e}^{-m/2}(1-\mathsf{e}^{-m/(2n)})\bigr)$. This is maximal for $n=1$ so we adopt $2m^{-1}(\mathsf{e}^{-m/2}-\mathsf{e}^{-m})$ as general upper bound valid for all positive integers $n$ and positive real numbers $m$. Regarding $S_{2}$, we use the inequality (21), and after the change $k\to n-k$, and then extending the range of $k$ to infinity and letting $\tau=n^{-1}m>0$, we get
$S_{2}\leq\sum_{1\leq k\leq n/2}mn^{-3}k^{2}\mathsf{e}^{-n^{-1}mk}\leq m^{-2}\sum_{k=1}^{\infty}\tau^{3}k^{2}\mathsf{e}^{-\tau k}=m^{-2}\tau^{3}\frac{\mathsf{e}^{-\tau}+\mathsf{e}^{-2\tau}}{(1-\mathsf{e}^{-\tau})^{3}}\,.$
It turns out that this function of $\tau>0$ is nicely decreasing with its supremum attained for $\tau\to 0^{+}$ (this can be proven using the Poisson summation formula). So $m^{2}S_{2}<2$. Adding the bound for $S_{1}$ we obtain (23) with
$K=2+\sup_{x>0}2x(\mathsf{e}^{-x/2}-\mathsf{e}^{-x}).$ (24)
The supremum is attained at $x_{0}\approx\numprint{2.89115}$ (and values decrease for $x>x_{0}$) and is $\approx\numprint{1.04138}$. So we can take $K=\numprint{3.05}$. Restricting to $m\geq m_{0}\geq 3$, we get (23) with $K=2+2m_{0}(\mathsf{e}^{-m_{0}/2}-\mathsf{e}^{-m_{0}})$, and only $2+m_{0}(\mathsf{e}^{-m_{0}/2}-\mathsf{e}^{-m_{0}})$ for even $n$’s. ∎
As a corollary, if we let both $m$ and $n$ go to infinity such that their ratio converges to some value $q$, we get (with $1$ on the right if $q=0$):
$\begin{split}\lim_{m,n\to\infty,mn^{-1}\to q}\frac{B_{m+1}(n)-B_{m+1}(0)}{n^{m+1}}=\lim\,(m+1)p_{m}(n)\\ =\lim\,mq_{m}(n)=\lim\frac{n^{-1}m}{\mathsf{e}^{n^{-1}m}-1}=\frac{q}{\mathsf{e}^{q}-1}\,.\end{split}$
This is indeed what one would expect from expressing the Bernoulli polynomials in terms of Bernoulli numbers and taking the limit term-wise, but Lemma 1 makes it easy (and gives an $O(m^{-1})$ error estimate). Reference [15], which uses the saddle-point method in the complex domain, provides full asymptotic for, among others, $B_{n}(nz+\frac{1}{2})$ as $n\to\infty$.
Lemma 2.
There is a constant $K_{2}$ such that for every positive $m$:
$0\leq\sum_{l=0}^{\lfloor\log_{2}(m+1)\rfloor}\Bigl((1-\mathsf{e}^{-m})q_{m}(2^{l})-p_{m}(2^{l})\Bigr)\leq\frac{K_{2}}{(m+1)^{2}}\,.$
Proof.
Let $L_{m}=\lfloor\log_{2}(m+1)\rfloor$. The $l=0$ term on the left-hand side of Equation (23) is, for $m>0$, $(1-\mathsf{e}^{-m})q_{m}(1)=(1-\mathsf{e}^{-m})(e^{m}-1)^{-1}=\mathsf{e}^{-m}\leq 4\mathsf{e}^{-1}(m+1)^{-2}$.
We assume now $m\geq 1$, so that $L_{m}\geq 1$, and we need only consider $1\leq l\leq L_{m}$. Using the same notation as in the proof of Lemma 1, and Equation (21), we have, for any even positive integer $n$:
$0\leq(1-\mathsf{e}^{-m})q_{m}(n)-p_{m}(n)\leq m^{-1}\mathsf{e}^{-m/2}+m^{-2}\sum_{k=1}^{\infty}\tau^{3}k^{2}\mathsf{e}^{-\tau k}\,$,
with $\tau=n^{-1}m$.
Summing for $2\leq 2^{l}\leq m+1$, the first term in the right-hand side contributes $L_{m}m^{-1}\mathsf{e}^{-m/2}$ which is a fortiori $O((m+1)^{-2})$. As per the second summands, as $\tau\geq m/(m+1)\geq\frac{1}{2}$, there holds $\mathsf{e}^{-\tau}\leq\mathsf{e}^{-\frac{1}{2}}$ and $\sum_{k\geq 1}k^{2}\mathsf{e}^{-\tau k}\leq\mathsf{e}^{-\tau}\sum_{k\geq 1}k^{2}\mathsf{e}^{-(k-1)/2}$. Hence, the second summand is, for each of $\tau=\frac{m}{2}$, $\frac{m}{4}$, …, $2^{-L_{m}}\geq\frac{1}{2}$, bounded by $Cm^{-2}\tau^{3}\mathsf{e}^{-\tau}$, for some constant $C$. We use the weaker upper bound $m^{-2}O(\mathsf{e}^{-\tau/2})$ and obtain in total $m^{-2}O(\mathsf{e}^{-1/4}+\mathsf{e}^{-1/2}+\dots+\mathsf{e}^{-2^{L_{m}}/8})=O(m^{-2})$. ∎
Let us set from here on, as in the previous proof:
$L_{m}=\lfloor\log_{2}(m+1)\rfloor\,.$
We also need an upper bound for $m\sum_{l=0}^{L_{m}}q_{m}(2^{l})=\sum_{l=0}^{L_{m}}\frac{m}{2^{l}(\mathsf{e}^{2^{-l}m}-1)}$. Not aiming at optimal results, we use the trivial bound valid for positive $x$:
$\frac{x}{\mathsf{e}^{x}-1}=\frac{x\mathsf{e}^{-x/2}}{2\sinh\frac{x}{2}}<\mathsf{e}^{-x/2}.$
Hence, as in the proof of Lemma 2, we get, supposing $m\geq 1$, hence $2^{L_{m}}\leq 2m$:
$m\sum_{l=0}^{L_{m}}q_{m}(2^{l})$ $\leq\mathsf{e}^{-\frac{m}{2}}+\mathsf{e}^{-\frac{m}{4}}+\dots+\mathsf{e}^{-2^{-L_{m}}\frac{m}{2}}$
$\leq\mathsf{e}^{-2^{L_{m}}/4}+\mathsf{e}^{-2^{L_{m}}/8}+\dots+\mathsf{e}^{-1/4}=O(1).$
For $0<m<1$, there is only one term $\frac{m}{\mathsf{e}^{m}-1}=O(1)$. Hence:
Lemma 3.
There is a constant $K_{3}$ such that for every positive $m$:
$0\leq\sum_{l=0}^{L_{m}}q_{m}(2^{l})\leq\frac{K_{3}}{m}\,.$
Lemma 4.
There is a constant $K_{4}$ such that for every positive $m$:
$0\leq\sum_{l=0}^{\lfloor\log_{2}(m+1)\rfloor}\Bigl(q_{m}(2^{l})-p_{m}(2^{l})\Bigr)\leq\frac{K_{4}}{m^{2}}\,.$
Proof.
From Lemmas 2 and 3 the sum is bounded above by $K_{2}(m+1)^{-2}+\mathsf{e}^{-m}K_{3}m^{-1}$, so, multiplying by $m^{2}$ it is bounded by $K_{2}+\mathsf{e}^{-1}K_{3}$. ∎
We now obtain:
Lemma 5.
There is a constant $K_{5}$ such that for every positive $m$:
$\sum_{l=0}^{L_{m}}\Bigr|mq_{m}(2^{l})-(m+1)p_{m}(2^{l})\Bigl|\leq\frac{K_{5}}{m+1}$
Proof.
For $0<m<1$, we have only one term $mq_{m}(1)-(m+1)p_{m}(1)$ with $mq_{m}(1)=m(e^{m}-1)^{-1}\leq 1$, and $p_{m}(1)=0$, so it is bounded by 11 hence by $K(m+1)^{-1}$ for any $K\geq 2$. For $m\geq 1$, we first decompose
$mq_{m}(2^{l})-(m+1)p_{m}(2^{l})=\begin{aligned} &(m+1)\Bigl((1-e^{-m})q_{m}(2^{l})-p_{m}(2^{l})\Bigr)\\ &-\bigl(1-(m+1)\mathsf{e}^{-m}\bigr)q_{m}(2^{l})\end{aligned},$
hence,
$\Bigl|mq_{m}(2^{l})-(m+1)p_{m}(2^{l})\Bigr|\leq(m+1)\Bigl((1-e^{-m})q_{m}(2^{l})-p_{m}(2^{l})\Bigr)+q_{m}(2^{l})\,.$
Using Lemmas 2 and 3 we get, recalling $m\geq 1$,
$\sum_{l=0}^{L_{m}}\Bigr|mq_{m}(2^{l})-(m+1)p_{m}(2^{l})\Bigl|\leq\frac{K_{2}}{m+1}+\frac{K_{3}}{m}\leq\frac{K_{2}+2K_{3}}{m+1}\,$,
and the statement to be proven holds with $K_{5}=\max(2,K_{2}+2K_{3})$. ∎
We now need to control the $2^{l}>m+1$ range. This is done via a completely different approach to the quantities $p_{m}(n)$ and $q_{m}(n)$ (or $q_{m+1}(n)$).
Lemma 6.
There are constants $K_{6}$ and $K_{7}$ such that
$\sum_{l>L_{m}}\Bigr|mq_{m}(2^{l})-(m+1)p_{m}(2^{l})\Bigl|$ $\leq\frac{K_{6}}{m+1}$ (25)
$\sum_{l>L_{m}}\Bigr|(m+1)q_{m+1}(2^{l})-(m+1)p_{m}(2^{l})\Bigl|$ $\leq\frac{K_{7}}{m+1}$ (26)
@@ 5. Approximations
In a certain way, Equation (26) is the one arising more naturally from the proof. But as we have already written up a proof for Lemma 5, which is expressed with the $q_{m}(2^{l})$’s, it is convenient to also prove Equation (25).
Proof.
Let $n$ be some positive integer. From Equation (20):
$(m+1)p_{m}(n)=\frac{B_{m+1}(n)-B_{m+1}(0)}{n^{m+1}}=\sum_{k=0}^{m}\binom{m+1}{k}B_{k}n^{-k}\,.$
We now approximate the binomial coefficients:
$\binom{m+1}{k}$ $=\frac{1}{k!}(m+1)^{k}\prod_{0\leq p<k}(1-(m+1)^{-1}p)$
$\implies 0$ $\leq\frac{(m+1)^{k}}{k!}-\binom{m+1}{k}\leq\frac{k(k-1)}{2(m+1)}\frac{(m+1)^{k}}{k!}\,.$
We used $\prod_{0\leq p<k}(1-\epsilon_{p})\geq 1-\sum_{0\leq p\leq k}\epsilon_{p}$ which is valid for $0\leq\epsilon_{p}\leq 1$, $0\leq p<k$. Hence
$\Bigl|\sum_{k=0}^{m}\frac{B_{k}}{k!}(m+1)^{k}n^{-k}-(m+1)p_{m}(n)\Bigr|\leq(m+1)^{-1}\sum_{k=0}^{m}\frac{|B_{k}|}{k!}\frac{k(k-1)}{2}\bigl((m+1)n^{-1}\bigr)^{k}$
We assume from here on $n\geq m+1$, which is the case for the powers of $2$ we will consider, $n=2^{l}$, $l>L_{m}$. The finite sums in the previous inequality are then partial sums of convergent series. We replace k(k1)k(k-1) by the larger $k(k+1)$ which verifies $m+1\leq\frac{k(k+1)}{2}$ for $k\geq m+1$, $m>0$. This allows to extend our termwise estimates to the $k>m$ range. Thus, there holds:
$\Bigl|\sum_{k=0}^{\infty}\frac{B_{k}}{k!}(m+1)^{k}n^{-k}-(m+1)p_{m}(n)\Bigr|\leq(m+1)^{-1}\sum_{k=2}^{\infty}\frac{|B_{k}|}{k!}\frac{k(k+1)}{2}\bigl((m+1)n^{-1}\bigr)^{k}.$
We recognize in the left-hand side
$\sum_{k=0}^{\infty}\frac{B_{k}}{k!}(m+1)^{k}n^{-k}=\frac{n^{-1}(m+1)}{\mathsf{e}^{n^{-1}(m+1)}-1}=(m+1)q_{m+1}(n)\,.$
The series $S(\tau)=\sum_{k\geq 2}^{\infty}\frac{|B_{k}|}{k!}\frac{k(k+1)}{2}\tau^{k}$ has radius of convergence $2\pi$, so there is some constant $C$ such that $S(\tau)\leq C_{1}\tau^{2}$ for $0\leq\tau\leq 1$. Hence:
$n\geq m+1\implies\Bigl|(m+1)q_{m+1}(n)-(m+1)p_{m}(n)\Bigr|\leq C_{1}(m+1)^{-1}\bigl((m+1)n^{-1}\bigr)^{2}.$
For $n=2^{l}$, $l>L_{m}$, we have $n>(m+1)$ indeed. And summing $\bigl((m+1)n^{-1}\bigr)^{2}$ over such powers of $2$ is bounded above by $1+4^{-1}+16^{-1}+\dots$. Hence Equation (26).
A modification in the handling the binomial coefficients will readily give Equation (25). We write this time:
$\binom{m+1}{k}=\frac{1}{k!}m^{k}\prod_{-1\leq p<k-1}(1-m^{-1}p)\,.$
The product (depending on $k\in[0,m]\cap\mathbb{Z}$) is bounded above by $1+m^{-1}$ and below by $\prod_{0\leq p<k-1}(1-m^{-1}p)$, hence the distance to $1$ is bounded in absolute value by
$m^{-1}\delta_{k},\qquad\text{with }\delta_{k}=\max(1,\frac{(k-1)(k-2)}{2})\qquad(k\geq 1).$
Moreover, $\delta_{0}=0$. We can use a weaker upper bound $k(k+1)/2$, so that $k(k+1)/2\geq m$ holds for $k\geq m+1$. This avoids having to discuss separately the (exponentially small, as one can check) contribution from the $k>m$ range. The sole differences with our earlier situation is that $k$ now starts at $1$ not at $2$ and that powers of $m+1$ are replaced by powers of $m$. And we obtain Equation (25). ∎
@@
Let, for positive real $m$,
$f_{m}=\sum_{l=0}^{\infty}\bigl(1-\frac{m2^{-l}}{\mathsf{e}^{m2^{-l}}-1}\bigr)\,.$ (27)
There holds $|e_{m}-f_{m}|\leq\frac{K}{m+1}$ for all $m>0$ and some constant $K$.